# The Largest Circle Enclosing $n$ Lattice Points

### v1

@@
Abstract. In this paper, we propose a class of elementary plane geometry problems closely related to the title of this paper. Here, a circle is the 1-dimensional curve bounding a disk. For any nonnegative integer, a circle is called $n$-enclosing if it contains exactly $n$ lattice points on the $xy$-plane in its interior. The main questions are when the largest $n$-enclosing circle exists and what the largest radius is. We study the small integer cases by hand and extend to all $n<1100$ with the aid of a computer. We find that frequently such a circle does not exist, e.g., when $n=5,6$. We then show a few general results on these circles including some regularities among their radii and an easy criterion to determine exactly when largest $n$-enclosing circles exist. Further, from numerical evidence, we conjecture that the set of integers whose largest enclosing circles exist is infinite, and so is its complementary in the set of nonnegative integers. Throughout this paper we present more mysteries/problems/conjectures than answers/solutions/theorems. In particular, we list many conjectures and some unsolved problems including possible higher dimensional generalizations at the end of the last two sections.
@@ 1 Introduction
In fact, the original problem did not have the condition on the area. Honsberger [6, p. 118] further strengthened it by restricting the area to be $n$. Several years earlier, Schinzel [12] showed that for every positive integer $n$, there exists a circle on the plane having exactly $n$ lattice points on its circumference, which was given the name Schinzel circle. Also see [6, p. 118]. For any circle of radius $r$ centered at the origin, it is well known that the number $N(r)$ of interior lattice points is given by
$N(r)=1+4\lfloor r\rfloor+\sum_{j=1}^{\lfloor r\rfloor}\lfloor\sqrt{r^{2}-j^{2}}\rfloor.$
See, e.g., [5]. This number is close to $\pi r^{2}$ for large $r$. In particular, one may find the following estimate credited to Gauss in Hardy’s book [4, p. 67]:
@@
We will present a related result in Theorem 6.1 with a short proof. Incidentally, the error bound for $E(r)$ can be further improved which turns out to be a significant problem having close connection with analytic number theory and therefore has attracted quite some work throughout the past century. See [2, 3, 7, 8, 10].
@@ 5.1 First special class
$f(k)\sim\pi R_{f(k)}^{2}=\pi\left(k^{2}+k+\frac{1}{2}\right).$
@@ 5.2 Second special class
$k$ 1 2 3 4 5 6 7 8 9
$g(k)$ 5 13 29 49 81 113 149 197 253
$\pi R_{g(k)}^{2}\approx$ 6 16 31 53 82 116 157 204 258
$k$ 10 11 12 13 14 15 16 17 18
$g(k)$ 317 377 441 529 613 709 797 901 1009
$\pi R_{g(k)}^{2}\approx$ 317 383 456 534 619 710 807 911 1021
Among all the $g(k)$’s in Table 3 the following are all strong MC: $k=4,7,8,11,12,16,17,18$. We will present their MC-circle symmetry in Figure 22 on page 22.
@@ 6 Complete classification of $n\leq 1000$
When $n$ gets larger and larger, it becomes more and more difficult to find the exact values of $R_{n}$. However, we have the following theoretical results to help us delegate the main computation to an algebra system such MAPLE or Mathematica.
@@
    $\sharp\left\{n\left|\begin{aligned} &{\mathcal{M}}_{n}\ \text{with only one }\\ &\text{hor./ver. symmetry}\end{aligned}\right.\right\}$ 22 21 17 24 22 23 23 21 26 12 17
@@
    $\sharp\left\{n\left|\begin{aligned} &{\mathcal{M}}_{n}\ \text{with at least }\\ &\text{two symmetries}\end{aligned}\right.\right\}$ 24 16 11 13 9 8 9 7 11 9 7
@@ 7 Concluding remarks and further research problems
The radius sequence $\{\rho_{n}\}_{n\geq 0}$ is not always increasing. However, whenever a drop after $\rho_{n}$ occurs there are some associated non-MC numbers. Moreover, the largest strictly increasing subsequence is conjectured to correspond exactly to all the MC numbers while the LUBOR $R_{n}$ of any non-MC $n$ is proved to be given by the largest radius in the sequence before $\rho_{n}$.
@@
    More generally, as Steinhaus’ Theorem 1.1 has been generalized to any Hilbert space $X$ in [9, 16], we wonder when the largest $n$-enclosing open balls exist for a quasi-finite set (see [16, p. 129] for the definition). Do we have similar results in this setting?
@@
MC number: maximally circlable number, i.e., a nonnegative number $n$ such that there is a a largest circle to contain $n$ lattice points in the interior
@@
MC-set of $n$: a set of $n$ latticed points that are enclosed by an MC-circle of $n$

### v2

@@
Abstract. In this paper, we propose a class of elementary plane geometry problems closely related to the title of this paper. Here, a circle is the 1-dimensional curve bounding a disk. For any nonnegative integer, a circle is called $n$-enclosing if it contains exactly $n$ lattice points on the $xy$-plane in its interior. The main questions are when the largest $n$-enclosing circle exists and what the largest radius is. We study the small integer cases by hand and extend to all $n<1100$ with the aid of a computer. We find that frequently such a circle does not exist, e.g., when $n=5,6$. We then show a few general results on these circles including some regularities among their radii and an easy criterion to determine exactly when the largest $n$-enclosing circles exist. Further, from numerical evidence, we conjecture that the set of integers whose largest enclosing circles exist is infinite, and so is its complementary in the set of nonnegative integers. Throughout this paper we present more mysteries/problems/conjectures than answers/solutions/theorems. In particular, we list many conjectures and some unsolved problems including possible higher dimensional generalizations at the end of the last two sections.
AMS Subject Classifications (2020): 152C05, 52C35, 52C15, 52C25.
@@ 1 Introduction
In fact, the original problem did not have the condition on the area. Honsberger [7, p. 118] further strengthened it by restricting the area to be $n$. Several years earlier, Schinzel [14] showed that for every positive integer $n$, there exists a circle on the plane having exactly $n$ lattice points on its circumference, which was given the name Schinzel circle. Also see [7, p. 118]. For any circle of radius $r$ centered at the origin, it is well known that the number $N(r)$ of lattice points in the interior and on the boundary is given by
$N(r)=1+4\lfloor r\rfloor+4\sum_{j=1}^{\lfloor r\rfloor}\lfloor\sqrt{r^{2}-j^{2}}\rfloor.$ (1)
See, e.g., [6] or [4, Eq. (8)]. The sequence $\{N(k)\}_{k\geq 1}$ is the sequence [A000328] on the OEIS website [17]. By considering the area one can see that the number $N(r)$ is close to $\pi r^{2}$ for large $r$. In particular, one may find the following estimate credited to Gauss in Hardy’s book [5, p. 67]:
@@
We will present a related result in Theorem 6.2 with a short proof. Incidentally, the error bound for $E(r)$ can be further improved which turns out to be a significant problem having close connection with analytic number theory and therefore has attracted quite some work throughout the past century. See [2, 3, 4, 8, 9, 10, 12].
@@ 5.1 First special class
$f(k)=4\sum_{j=1}^{k}\left(\left\lceil\sqrt{k^{2}+k+\frac{1}{2}-\Big{(}j-\frac{1}{2}\Big{)}^{2}}+\frac{1}{2}\right\rceil-1\right)\sim\pi R_{f(k)}^{2}=\pi\left(k^{2}+k+\frac{1}{2}\right).$
@@ 5.2 Second special class
$k$ 1 2 3 4 5 6 7 8 9 10
$g(k)$ 5 13 29 49 81 113 149 197 253 317
$\pi R_{g(k)}^{2}\approx$ 6 16 31 53 82 116 157 204 258 317
$k$ 11 12 13 14 15 16 17 18 19 20
$g(k)$ 377 441 529 613 709 797 901 1009 1129 1257
$\pi R_{g(k)}^{2}\approx$ 383 456 534 619 710 807 911 1021 1137 1260
Among all the $g(k)$’s in Table 3 with $g(k)<1100$ the following are all strong MC: $k=4,7,8,11,12,16,17,18$. We will present their MC-circle symmetry in Figure 22 on page 22.
Notice that the sequence $\{g(k)\}_{k\geq 1}$ is [A000328] on the OEIS website [17]. This is equivalent to saying that when we shrink the circle ${\mathcal{T}}_{k}$ to radius $k$, no interior points are moved outside. Indeed, if lattice point $(a,b)$ satisfies $a^{2}+b^{2}<k^{2}+1$ if and only if $a^{2}+b^{2}\leq k^{2}$.
@@ 6 Complete classification of $n\leq 1000$
When $n$ gets larger and larger, it becomes more and more difficult to find the exact values of $R_{n}$. However, we have the following theoretical results to help us delegate the main computation to an algebra system such MAPLE or Mathematica. The first one is an analogous formula of (1) for the number of interior points inside a circle of radius $r$.
Theorem 6.1.
For any circle of radius $r$ centered at the origin, the number $\nu(r)$ of interior lattice points is given by
$\nu(r)=4\lceil r\rceil-3+4\sum_{j=1}^{\lceil r\rceil-1}(\lceil\sqrt{r^{2}-j^{2}}\rceil-1).$ (3)
Proof.
It is clear that the number of interior lattice points on the $x$- or $y$-axes is
$31+4(\lceil r\rceil-1)=4\lceil r\rceil-$ 3
since $\lceil r\rceil=r$ if $r$ is an integer while $\lceil r\rceil-1$ is the integer part of $r$ otherwise. Suppose now $(j,k)$ is in the interior of the circle with $j,k>0$. Then they must satisfy $11\leq j\leq\lceil r\rceil-$ 1 and $j^{2}+k^{2}<r^{2}$, whence the summands in (3). ∎
@@
Remark 6.3.
The sequence $\{\nu(k)\}_{k\geq 1}$ is the sequence A051132 on the OEIS website [17]. The difference sequence $\{N(k)-\nu(k)\}_{k\geq 1}$ [A046109] provides the number of lattice points on the circumference of the circle with radius $r$ centered at 00, where $N(r)$ is defined by (1).
@@
  3. (3)
    Let $n\neq 6$ and let $\rho_{n}$ be the radius of the largest $n$-enclosing lattice circle if it exists and $\rho_{n}=0$ otherwise. Then the numbers in $\{\rho_{n}\}_{n\geq 0}$ are pairwise distinct.
@@
  3. (3)
    This holds for all $11110\leq n<$ $1$. In particular, it implies the proceeding conjecture which is equivalent to saying $\rho_{n}\neq 0$ unless $n=6$.
@@
    $\sharp\left\{n\left|\begin{aligned} &{\mathcal{M}}_{n}\ \text{has only one }\\ &\text{hor./ver. symmetry}\end{aligned}\right.\right\}$ 22 21 17 24 22 23 23 21 26 12 17
@@
    $\sharp\left\{n\left|\begin{aligned} &{\mathcal{M}}_{n}\ \text{has at least }\\ &\text{two symmetries}\end{aligned}\right.\right\}$ 24 16 11 13 9 8 9 7 11 9 7
@@ 7 Concluding remarks and further research problems
The radius sequence $\{\rho_{n}\}_{n\geq 0}$ is not always increasing. However, whenever a drop after $\rho_{n}$ occurs there are some associated non-MC numbers. Moreover, the largest strictly increasing subsequence is conjectured to correspond exactly to all the MC numbers while the LUBOR $R_{n}$ of any non-MC $n$ is proved to be given by the largest radius in the sequence before $\rho_{n}$. When we only consider the largest lattice $n$-enclosing circles, $n=129$ is the only number less than 1100 such that $\rho_{n}<\rho_{n-1}$ with both $n$ and $n-1$ are non MC numbers.
@@
    More generally, as Steinhaus’ Theorem 1.1 has been generalized to any Hilbert space $X$ in [11, 19], we wonder when the largest $n$-enclosing open balls exist for a quasi-finite set (see [19, p. 129] for the definition). Do we have similar results in this setting?
@@
MC number: maximally circlable number, i.e., a nonnegative number $n$ such that there is a largest circle to contain $n$ lattice points in the interior
@@
MC-set of $n$: a set of $n$ lattice points that are enclosed by an MC-circle of $n$